A freshly caught catfish is placed on a spring scale, and it oscillates up and down with a period of 0.2 s.

If the spring constant of the scale is 2160 N/m, what is the mass of the catfish?
Answer #1

Solution)

We know, Time Period T= 2*pi*sqrt(m/k)

So, mass m= T^2*k/4*pi^2

Mass, m= 0.2^2 * 2160/4*pi^2 = 2.190 kg (Ans)

=======

Good luck!:)